\newcommand{\etalchar}[1]{$^{#1}$}
\begin{thebibliography}{AHS{\etalchar{+}}24}

\bibitem[AHS{\etalchar{+}}24]{Atalig}
Sunny Atalig, Alexander Hickerson, Arrdya Srivastav, Tingting Zheng, and Marek Chrobak.
\newblock Lower bounds for adaptive relaxation-based algorithms for single-source shortest paths.
\newblock In {\em {ISAAC}}, volume 322 of {\em LIPIcs}, pages 8:1--8:16. Schloss Dagstuhl - Leibniz-Zentrum f{\"{u}}r Informatik, 2024.

\bibitem[BE12]{Bannister}
Michael~J. Bannister and David Eppstein.
\newblock Randomized speedup of the {B}ellman--{F}ord algorithm.
\newblock In {\em Proceedings of the Ninth Workshop on Analytic Algorithmics and Combinatorics (ANALCO)}, pages 41--47. SIAM, 2012.

\bibitem[Bel58]{Bellman}
Richard Bellman.
\newblock On a routing problem.
\newblock {\em Quarterly of applied mathematics}, 16(1):87--90, 1958.

\bibitem[Epp23]{Eppstein}
David Eppstein.
\newblock Lower bounds for non-adaptive shortest path relaxation.
\newblock In {\em Algorithms and Data Structures - 18th International Symposium, {WADS} 2023, Proceedings}, volume 14079 of {\em Lecture Notes in Computer Science}, pages 416--429. Springer, 2023.

\bibitem[EV20]{engen2020containing}
Michael Engen and Vincent Vatter.
\newblock Containing all permutations.
\newblock {\em The American Mathematical Monthly}, 128(1):4--24, 2020.

\bibitem[Fin24]{Fineman}
Jeremy~T. Fineman.
\newblock Single-source shortest paths with negative real weights in \emph{{\~{O}}(mn\({}^{\mbox{8/9}}\))} time.
\newblock In {\em Proceedings of the 56th Annual {ACM} Symposium on Theory of Computing, {STOC} 2024, Vancouver, BC, Canada, June 24-28, 2024}, pages 3--14. {ACM}, 2024.

\bibitem[For56]{Ford}
Lester~Randolph Ford.
\newblock Network flow theory.
\newblock 1956.

\bibitem[Hed88]{rfc1058}
Charles Hedrick.
\newblock {\em {Routing Information Protocol}}, June 1988.
\newblock RFC 1058.

\bibitem[HJQ25]{huang2025faster}
Yufan Huang, Peter Jin, and Kent Quanrud.
\newblock Faster single-source shortest paths with negative real weights via proper hop distance.
\newblock In {\em Proceedings of the 2025 Annual ACM-SIAM Symposium on Discrete Algorithms (SODA)}, pages 5239--5244. SIAM, 2025.

\bibitem[HS88]{haddad1988recognizing}
Ramsey~W. Haddad and Alejandro~A. Sch{\"a}ffer.
\newblock Recognizing {B}ellman--{F}ord-orderable graphs.
\newblock {\em SIAM journal on discrete mathematics}, 1(4):447--471, 1988.

\bibitem[KK76]{kleitman}
Daniel~J. Kleitman and David~Joseph Kwiatkowski.
\newblock A lower bound on the length of a sequence containing all permutations as subsequences.
\newblock {\em Journal of Combinatorial Theory, Series A}, 21(2):129--136, 1976.

\bibitem[Mai78]{superseq1}
David Maier.
\newblock The complexity of some problems on subsequences and supersequences.
\newblock {\em J. {ACM}}, 25(2):322--336, 1978.

\bibitem[Moo59]{Moore}
Edward~F. Moore.
\newblock The shortest path through a maze.
\newblock In {\em Proc. of the International Symposium on the Theory of Switching}, pages 285--292. Harvard University Press, 1959.

\bibitem[MR85]{mendelsohn1985one}
Eric Mendelsohn and Alexander Rosa.
\newblock One-factorizations of the complete graph—a survey.
\newblock {\em Journal of Graph Theory}, 9(1):43--65, 1985.

\bibitem[MR17]{routing}
Deep Medhi and Karthik Ramasamy.
\newblock {\em Network routing: algorithms, protocols, and architectures}.
\newblock Morgan Kaufmann, 2017.

\bibitem[MS86]{Mehlhorn_BF}
Kurt Mehlhorn and Bernd~H. Schmidt.
\newblock On {BF}-orderable graphs.
\newblock {\em Discrete applied mathematics}, 15(2-3):315--327, 1986.

\bibitem[Rad12]{Radomirovic}
Sasa Radomirovic.
\newblock A construction of short sequences containing all permutations of a set as subsequences.
\newblock {\em the electronic journal of combinatorics}, 19(4):P31, 2012.

\bibitem[RU81]{superseq2}
Kari{-}Jouko R{\"{a}}ih{\"{a}} and Esko Ukkonen.
\newblock The shortest common supersequence problem over binary alphabet is {NP}-complete.
\newblock {\em Theor. Comput. Sci.}, 16:187--198, 1981.

\bibitem[Uzn15]{Uznanski}
Przemys{\l}aw Uzna{\'n}ski.
\newblock All permutations supersequence is {coNP}-complete.
\newblock {\em arXiv preprint arXiv:1506.05079}, 2015.

\bibitem[Yen70]{Yen}
Jin~Y. Yen.
\newblock An algorithm for finding shortest routes from all source nodes to a given destination in general networks.
\newblock {\em Quarterly of applied mathematics}, 27(4):526--530, 1970.

\end{thebibliography}

\documentclass[abstract=true, DIV=14, parskip=half]{scrartcl}

\usepackage[noEnd,commentColor=black]{algpseudocodex}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{amsthm}
\usepackage[mathscr]{euscript}
\usepackage[shortlabels]{enumitem}
\usepackage{graphicx} %
\usepackage{subcaption}
\usepackage{thmtools}
\usepackage{todonotes}
\usepackage{xcolor}
\usepackage{microtype}

\makeatletter
\def\@fnsymbol#1{\ensuremath{\ifcase#1\or \!\;\or \!\;\or \ddagger\or
   \mathsection\or \mathparagraph\or \|\or **\or \dagger\dagger
   \or \ddagger\ddagger \else\@ctrerr\fi}}
\makeatother

\usepackage[normalem]{ulem}
\usepackage{tcolorbox}

\usepackage{upgreek}
\usepackage[colorlinks]{hyperref}
\usepackage[capitalize]{cleveref}

\crefname{equation}{eq.}{eqs.} %
\crefname{enumi}{}{} %
\crefname{icase}{case}{cases}
\crefname{ipart}{part}{parts}
\crefname{iprop}{property}{properties}
\crefname{iinv}{invariant}{invariants}

\crefformat{section}{\S\,#2#1#3}
\crefformat{subsection}{\S\,#2#1#3}
\crefrangeformat{section}{\S\S\,#3#1#4--#5#2#6}
\crefmultiformat{section}{\S\S\,#2#1#3}{,~#2#1#3}{,~#2#1#3}{,~#2#1#3}

\usepackage{authblk}
\usepackage[normalem]{ulem}

\usepackage{tikz}
\usetikzlibrary{calc}

\DeclareMathOperator{\relx}{\mathsf{relax}}

\renewcommand{\alpha}{\upalpha}

\newcommand{\ceil}[1]{\lceil #1 \rceil}

\newcommand{\connected}[1]{\def\temp{#1}\ifx\temp\empty\sim\else\overset{#1}{\sim}\fi}

\newtheorem{theorem}{Theorem}[section]
\newtheorem{lemma}[theorem]{Lemma}

\theoremstyle{definition}

\author{Jialu Hu}
\author{L\'{a}szl\'{o} Kozma}
\affil{Institut für Informatik, Freie Universit\"at Berlin, Germany \authorcr \texttt{firstname.lastname@fu-berlin.de}}

\title{Non-adaptive Bellman-Ford:\\ Yen's improvement is optimal\thanks{$^{*}$Supported by DFG Grant KO 6140/1-2.}}

\date{}

\begin{document}

\maketitle

\vspace{-0.1in}
\begin{abstract}
The Bellman-Ford algorithm for single-source shortest paths repeatedly updates tentative distances in an operation called \emph{relaxing an edge}. In several important applications a \emph{non-adaptive} (oblivious) implementation is preferred, which means fixing the entire sequence of relaxations upfront, independently of the edge-weights. Such an implementation performs, in a dense graph on $n$ vertices, $(1+o(1))n^3 $ relaxations. An improvement by Yen from 1970 reduces the number of relaxations by a factor of two. We show that no further constant-factor improvements are possible, and every \emph{non-adaptive deterministic} algorithm based on relaxations must perform $(\frac{1}{2} - o(1))n^3$ steps. This improves an earlier lower bound of Eppstein of $(\frac{1}{6} - o(1))n^3$. Given that a \emph{non-adaptive randomized} variant of Bellman-Ford with at most $(\frac{1}{3} + o(1))n^3$ relaxations (with high probability) is known, our result implies a strict separation between deterministic and randomized strategies, answering an open question of Eppstein.  
On the complexity side, we show that deciding whether a given relaxation sequence is guaranteed to yield correct distances is NP-hard, even with the complete graph as input. 
\end{abstract}

\section{Introduction}
The Bellman-Ford algorithm~\cite{Bellman, Ford, Moore} is a classical method for computing \emph{single-source shortest paths} in edge-weighted, directed graphs. Similarly to Dijkstra's algorithm, the Bellman-Ford algorithm uses \emph{edge relaxations}. Relaxing an edge $(u,v)$ means setting the tentative distance from the source $s$ to $v$ to the tentative distances from $s$ to $u$ plus the weight of the edge $(u,v)$, if this is smaller than the current tentative distance from $s$ to $v$. Initially setting the tentative distance from $s$ to $s$ itself to $0$ and to all other vertices to $+\infty$, the algorithm proceeds via repeated edge-relaxations, updating tentative distances until they reach their final values, the true distances. Relaxing an edge $(u,v)$ changes the tentative distance from $s$ to $v$ to its final (correct) value, if and only if the tentative distance from $s$ to $u$ is already correct and $(u,v)$ is the last edge of a shortest path from $s$ to $v$. Algorithms of this type are also known as \emph{label-correcting}.

Besides the correct handling of negative edge-weights, an important feature of the Bellman-Ford algorithm is that it can be executed in a \emph{non-adaptive} (also called \emph{oblivious}) way. This means that the sequence of edge-relaxations is fixed in advance, independently of the edge-weights of the input graph and of the outcomes of previous operations. This reduces overhead while maintaining the same asymptotic worst-case running time. A non-adaptive sequence of edge-relaxations also makes the algorithm amenable for implementation in parallel or distributed settings, e.g., in network routing protocols~\cite{rfc1058, routing}, where centralized control may not be available.  

It is easy to see that a sequence of (non-adaptive) edge-relaxations correctly computes the distance from $s$ to $v$ if and only if the edges of some shortest path from $s$ to $v$ are relaxed in the order in which they appear on the path (possibly with other edge-relaxations in between). If for every simple path from $s$ to $v$, the edges of the path are relaxed in the order they appear going from $s$ to $v$, then the correct distance from $s$ to $v$ must have been computed. Conversely, if the edges of some simple $s$-to-$v$ path do not appear as a subsequence of the relaxation sequence, then, for some configuration of the edge-weights (e.g., with weights on the path set to $0$ and all other weights set to $1$) the correct distance will not have been reached. 

The problem of finding the most efficient edge-relaxation sequence thus maps to the combinatorial problem of finding the shortest sequence of edges in which every simple path from $s$ to another vertex appears as a subsequence. The standard non-adaptive version of the Bellman-Ford algorithm relaxes all $m$  edges in a round-robin fashion $n-1$ times, where $n$ is the number of vertices. The resulting sequence of edges clearly contains every simple path, and in a complete directed graph the number of edge-relaxations is $(1+o(1))n^3$, with a proportional overall running time.    

A classical improvement by Yen~\cite{Yen} reduces the number of relaxations by a factor of two, while preserving the non-adaptive nature of the algorithm. Yen's algorithm decomposes the edge set of the input graph into \emph{forward edges} and \emph{backward edges}, both sets forming acyclic subgraphs. It then alternates between relaxing all forward- and all backward edges in topological order. 

The question of whether a better relaxation order exists than the one given by Yen's method has been open since 1970. Recently, Eppstein~\cite{Eppstein} showed that every non-adaptive deterministic algorithm must perform $(\frac{1}{6} - o(1))n^3$ relaxations in dense graphs, leaving open the possibility of a further factor-three improvement.
Bannister and Eppstein~\cite{Bannister} obtained a \emph{randomized} improvement to Yen's algorithm, with a high-probability guarantee on the algorithm's success. In the non-adaptive setting, this improves Yen's bound by a factor of $\frac{2}{3}$.

\paragraph{Our results.} We show (\S\,\ref{sec2}) that every non-adaptive deterministic algorithm based on edge-relaxations must perform $(\frac{1}{2} - o(1))n^3$ relaxation steps on a complete directed graph with $n$ vertices. This shows that Yen's algorithm cannot be improved with the choice of a better relaxation sequence, by any constant factor. Moreover, since faster randomized approaches are known, randomization strictly helps for non-adaptive relaxation sequences. This answers an open question of Eppstein~\cite{Eppstein}.

On the complexity side, two natural questions arise. Given a directed graph $G$ and a source vertex $s$, we may ask for a \emph{shortest} non-adaptive sequence of edge-relaxations that compute the correct distances from $s$, for all possible configurations of edge-weights. In light of the previous discussion, this amounts to finding a shortest sequence of edges that contains as subsequence every simple path from $s$ in $G$. A second question is to decide, whether a \emph{given} sequence $S$ of edges is a valid relaxation sequence, i.e., whether it contains every simple path from $s$ in $G$. 

While we suspect both problems to be (NP-)hard, we can only show this for the second, and leave the first one open. The hardness of the second problem holds even for the complete graph $G$ (\S\,\ref{sec3}). 

We note that both problems admit polynomial-time algorithms in the easier case of \emph{BF-orderable} graphs, i.e., graphs in which every edge needs to be relaxed at most once; this results from 1980s work of Mehlhorn and Schmidt~\cite{Mehlhorn_BF}, and Haddad and Sch\"{a}ffer~\cite{haddad1988recognizing}.

\paragraph{Further related work.}
Due to the fundamental nature of the single-source shortest paths problem in the presence of negative weights, several attempts were made to establish its precise complexity or to improve the long-standing $O(mn)$ runtime of the Bellman-Ford algorithm. We refer to Eppstein~\cite{Eppstein} for a discussion of results in different models and for variants of the problem. Note that complexity bounds are sensitive to the exact modeling assumptions, in particular, they differ between the non-adaptive and adaptive settings. 

When adaptivity is allowed, i.e., the algorithm can decide which edge to relax based on the outcomes of past relaxations (and possibly other calculations involving edge-weights), a recent breakthrough by Fineman~\cite{Fineman} has reduced the complexity to $\tilde{O}(mn^{8/9})$, with a later improvement~\cite{huang2025faster} to $\tilde{O}(mn^{4/5})$. These results may explain why \emph{lower bounds} for adaptive algorithms have been elusive. Nevertheless, an asymptotic lower bound of $\Omega(n^3)$ has recently been shown~\cite{Atalig}, for adaptive relaxation-based algorithms with a certain restricted set of operations (while classical Bellman-Ford-variants fall within these restrictions, the mentioned recent improvements do not). 

\section{Deterministic lower bound}\label{sec2}
Let $G$ be the complete directed graph with vertex set $V$ and edge set $E$, with $|V| = n$ and $|E| = n(n-1)$, and let $s \in V$ be the source vertex. Let $w:E \rightarrow \mathbb{R}$ denote the edge-weights. Note that $G$ is simple, containing no loops or multiple edges. 

The Bellman-Ford algorithm and its variants compute the distances from $s$ to every vertex in $V$ by initially setting $d(s,s) = 0$ and $d(s,v) = +\infty$ for all $v \in V \setminus \{s\}$, and repeatedly performing operations $\relx(e)$ for edges $e \in E$. The operation $\relx(e)$ for $e = (u,v)$ sets $d(s,v) \leftarrow \min\{d(s,v), d(s,u)+w(e)\}$. The goal is to execute such steps until $d(s,v)$ is the correct distance from $s$ to $v$ for all $v \in V$. We assume that the graph contains no negative weight cycle (such a cycle can be detected by checking if any relax operation still reduces tentative distances after the algorithm has finished). 

As we consider non-adaptive algorithms, the behaviour of the algorithm is fully specified by the sequence of relax operations $S = (e_1, \dots, e_L)$ it performs, where $e_i \in E$ for all $1 \leq i \leq L$. We refer to $L = |S|$ as the length of the sequence. Our main result is the following.

\begin{theorem}\label{thm1}
Every valid relaxation sequence for $G$ must have length at least $(\frac{1}{2}-o(1))n^3$.
\end{theorem}

In the remainder of the section we prove the theorem. All paths considered are simple. We say that a sequence $S$ of edges \emph{contains} a path $(v_1, \dots, v_k)$ if $S$ contains $(v_1,v_2), (v_2,v_3), \dots, (v_{k-1},v_k)$ as a subsequence (in the given order, but not necessarily contiguously). A \emph{$k$-path} is a path $(s,v_1,\dots,v_k)$ for some set of $k$ vertices $\{v_1, \dots, v_k\} \subseteq V \setminus \{s\}$. We say that a sequence of edges $S$ is \emph{$k$-valid}, if it contains every possible $k$-path. By our earlier discussion, a sequence $S$ is a valid relaxation sequence for $G$ in a non-adaptive single-source shortest path algorithm, if and only if it is $(n-1)$-valid.  

The task of finding a short $k$-valid sequence appears related to the task of finding a short \emph{permutation supersequence}, a well-studied question in combinatorics. 
In that context, $c_n$ denotes the smallest possible length of a sequence over $\{1,\dots,n\}$ that contains each permutation of $\{1,\dots,n\}$ as a subsequence (not necessarily contiguously).

It is known that $c_n \leq \ceil{n^2-\frac{7}{3}n + \frac{19}{3}}$~\cite{Radomirovic}, see also the survey of Engen and Vatter~\cite{engen2020containing}. Most relevantly for us, a lower bound of the form $c_n \geq (1- o(1))n^2$ was shown in 1976 by Kleitman and Kwiatkowski.

\begin{lemma}[\cite{kleitman}]\label{lem1}
For all $n \geq 1$ and all $\varepsilon > 0$, we have $c_n > n^2 - C_\varepsilon \cdot n^{7/4 + \varepsilon}$, with $C_\varepsilon$ depending only on $\varepsilon$. 
\end{lemma}

We will use Lemma~\ref{lem1} to prove Theorem~\ref{thm1}.\footnote{In an earlier version of the paper we gave a more direct proof of Theorem~\ref{thm1}, but the proof was faulty.}

It is well-known that whenever $N$ is even, the edge-set of a complete (undirected) graph on $N$ vertices can be decomposed into $N-1$ disjoint perfect matchings, i.e., sets of $\frac{N}{2}$ non-incident edges. (Such a decomposition is called a \emph{$1$-factorization}, often studied in the context of round-robin tournaments, e.g., see~\cite{mendelsohn1985one}). If $N$ is odd, we can simply ignore one of the vertices and decompose the remainder of the graph into $N-2$ disjoint matchings, each consisting of $\frac{N-1}2$ edges. 

Denote the vertex set of the complete directed graph $G$ as $V = \{s,v_1,\dots,v_{n-1}\}$. 
Consider a complete (undirected) graph on $\{v_1, \dots, v_{n-1}\}$. By the previous paragraph, regardless of the parity of $n$, we can find in this graph $n-3$ disjoint matchings, each of size $\frac{n-2}{2}$. Denote such a collection of matchings as $E_1, \dots, E_{n-3}$ (each $E_i$ is a set of undirected edges).

For each matching $E_i$ we form two sets of \emph{directed} edges in $G$; one where each edge of $E_i$ is oriented from smaller to larger index vertex (call this $F_i$), and one where each edge of $E_i$ is oriented from larger to smaller index vertex (call this ${H_i}$). We thus obtain a collection $\mathcal{F} = \{F_1,H_1, \dots, F_{n-3}, {H_{n-3}}\}$ of pairwise disjoint sets, each consisting of pairwise non-incident directed edges in $G$. 

Let $S$ be a shortest $(n-1)$-valid, or simply valid, sequence of edges for graph $G$. For any $F \in \mathcal{F}$ let $S_F$ denote the maximal subsequence of $S$ consisting of elements in $F$.

We claim that $S_F$ must contain every permutation of $F$ as a subsequence, and thus the lower bound of Lemma~\ref{lem1} applies to $S_F$.

Indeed, suppose some permutation of $F$ is not contained in $S_F$, and denote this permutation (of directed edges of $G$) as $(e_1, \dots, e_{\frac{n-2}{2}})$, where $e_i$ is the directed edge $(x_i, y_i)$, for $x_i,y_i \in V \setminus \{s\}$. 
In this case, $S$ could not contain the sequence of edges $( (s,x_1), \mathbf{e_1}, (y_1,x_2), \mathbf{e_2}, (y_2,x_3), \dots, \mathbf{e_{\frac{n-2}{2}}} )$ that form a simple path from the source to $y_{\frac{n-2}{2}}$, and thus $S$ could not be valid. 

Since $|F| = \frac{n-2}{2}$ for all $F \in \mathcal{F}$, we have $|S_F| \geq (\frac{1}{4}-o(1))n^2$, by Lemma~\ref{lem1}. As edge sets in $\mathcal{F}$ are pairwise disjoint, the corresponding subsequences of $S$ do not overlap. Thus, we have $$|S| ~\geq~ |\mathcal{F}| \cdot \left(\frac{1}{4}-o(1)\right)n^2 ~\geq~ \left(2n-6\right)\left(\frac{1}{4}-o(1)\right)n^2 ~\geq~ \left(\frac{1}{2}-o(1)\right)n^3.$$ This completes the proof. $\hfill \qed$

{} %

\section{Hardness of deciding validity}\label{sec3}

Again, let $G$ be a directed graph, with a special vertex $s$. 
Let $S = (e_1, \dots, e_{|S|})$ be a sequence of edges of $G$. The \emph{relaxation validity} question asks whether $S$ contains as subsequence every simple path of $G$ starting in $s$, in other words, whether $S$ is a valid relaxation sequence. %

\begin{theorem}\label{thm2}
    Relaxation validity is co-NP-complete.
\end{theorem}

In the remainder of the section we prove Theorem~\ref{thm2}. Membership in co-NP is trivial; if $S$ is not valid, then some simple path $P$ witnesses this. 
Verifying that $S$ does not contain $P$ can be done by greedily matching edges of $P$ in their order from $s$, to the earliest occurrence in $S$, to the right of the entry matched to the predecessor edge. 

To show co-NP-hardness, we reduce from the \emph{all-permutations supersequence} (APS) problem, shown to be co-NP-hard by Uzna\'{n}ski~\cite{Uznanski}.
APS asks whether a given sequence $T = (t_1, \dots, t_{|T|})$ contains every permutation of $\{1,\dots,n\}$ as a subsequence. (Alternatively, one may ask whether there is a permutation \emph{not} contained in $T$, which is consequently NP-hard).

The reduction from APS to relaxation validity is straightforward. Consider a complete directed graph $G$ with a set of $2n+1$ vertices denoted $V = \{s,x_1,y_1, \dots, x_n,y_n\}$. For $i \in \{1,\dots,n\}$, let $e_i$ denote the special edge $(x_i, y_i)$. 

Let $Q$ be a valid relaxation sequence of $G$, i.e., a $2n$-valid sequence of edges. Let $R$ be the sequence of edges obtained from $Q$ by deleting every occurrence of $e_i$ for all $i$.

We construct a relaxation sequence $S$ of $G$ as follows. For $i = 1,\dots,|T|$, we form $S_i$ by appending to the end of sequence $R$ the edge $e_j$, for $j=t_i$. Then, we concatenate $S_1, \dots, S_{|T|}$, and finally $R$, to form the sequence $S$. From the preceding discussion, we can enforce $|Q|,|R| \in O(n^3)$ and the construction of $S$ takes time polynomial in $|T|$. (Notice that we do not require $Q$ to be optimal, a simple round-robin sequence of length $O(n^3)$ is sufficient).

We claim that $T$ contains every permutation of $\{1, \dots, n\}$ if and only if $S$ contains every simple path starting from $s$ in $G$, from which Theorem~\ref{thm2} follows.

First, suppose $T$ does not contain some permutation $\pi_1, \dots, \pi_n$. Then, $S$ does not contain the sequence of edges $(e_{\pi_1}, \dots, e_{\pi_n})$ as these edges only appear in $S$ in correspondence to entries of $T$. Consequently, $S$ cannot correctly relax the path $(s,x_{\pi_1}, y_{\pi_1}, \dots, x_{\pi_n},y_{\pi_n})$.

Suppose now that $T$ contains every permutation of $\{1, \dots, n\}$. Then, every simple $k$-path $P$ is correctly relaxed by $S$. 
Indeed, since $T$ contains every permutation, $S$ contains the (possibly incomplete) permutation of edges $e_{\pi_1}, e_{\pi_2}, \dots$ in $P$, in their order of appearance in $P$. The portions of $P$ between any two edges $e_{\pi_i}$, $e_{\pi_{i+1}}$ or before the first, respectively, after the last such edge, are necessarily contained in some copy of $R$ inserted between the corresponding occurrences of $e_{\pi_i}$, $e_{\pi_{i+1}}$ in $S$, resp., at the beginning or at the end of $S$. This concludes the proof.   \hfill $\qed$
\section{Conclusions}

We showed that the number of edge-relaxations in Yen's variant of the Bellman-Ford algorithm cannot be improved in a deterministic non-adaptive way by any constant factor. Our lower bound is shown for the complete directed graph as the input. Besides serving as a worst-case example, the complete graph is also natural as it captures \emph{full non-adaptivity}: If an algorithm is oblivious to both the weights and the \emph{structure} of the graph, then it must attempt to relax every possible edge, as if the input were a complete graph. 

In the randomized case, a gap between the $(\frac{1}{12} - o(1))n^3$ lower bound due to Eppstein and the $(\frac{1}{3} + o(1))n^3$ upper bound due to Bannister and Eppstein remains~\cite{Eppstein, Bannister}.
For incomplete \emph{dense} graphs, i.e., with the number of edges $m \geq n^{1+ {\varepsilon}}$, Eppstein shows that in the worst case, the $O(mn)$ bound cannot be improved by non-adaptive relaxation sequences (in this case, the algorithm knows the graph structure but not the weights). For sparse families, a gap between the $O(mn)$ upper bound and an $\Omega(mn/\log{n})$ lower bound remains~\cite{Eppstein}. 

The complexity of finding a short relaxation sequence for a given graph remains open. The problem appears similar to the \emph{shortest supersequence} problem, known to be NP-hard in several variants (e.g., see~\cite{superseq1, superseq2}). The difficulty of extending these results to relaxation sequences is twofold. The paths in our setting are \emph{simple}, i.e., without repeated vertices or edges, whereas in the inputs to the shortest supersequence problem symbols can appear multiple times. To our knowledge, the complexity of finding the shortest supersequence of a set of (partial) permutations is open. A second difficulty is that the set of input paths in our setting is only implicitly given and the number of paths is, in general, exponential. 

\newpage

\small
\bibliographystyle{alpha}
\bibliography{main}

\end{document}